OpenStudy (anonymous):

The Green family and the Robinson family each used their sprinklers last summer. The water output rate for the Green family's sprinkler was 15L per hour. The water output rate for the Robinson family's sprinkler was 25L per hour. The families used their sprinklers for a combined total of 55 hours, resulting in a total water output of 1125L . How long was each sprinkler used?

OpenStudy (anonymous):

We note with: G-numbers of hours that Green family used sprinklers. R-numbers of hours that Robinson family used sprinklers. We have this equations: \[1*G+1*R=55\]\[15*G+25*R=1125\]=> \[G=55-R\]\[15(55-R)+25R=1125=>825-15R+25R=1125=>10R=300=>R=30\] =>\[G=25\]\[R=30\]